A fort has food for 160 people for 50 days . After 10 days 20 ppl left the fort . How long the food last at the same rate for the remaining ppl ?
Answer by ptaylor(2198)

A fort has food for 160 people for 50 days . After 10 days 20 ppl left the fort . How long the food last at the same rate for the remaining ppl ?
Let x= amount of time the food will last for the remaining people
The fort has 160*50=8000 people-days worth of food
In 10 days the fort consumes 160*10=1600 people-days of food leaving (8000-1600)=6400 people-days of food yet to be consumed by (160-20)=140 people
Sooooo, our equation to solve is:
140x=6400
x=45.714 days---amount of time the food will last for the remaining people
Hope this helps---ptaylor